Find all solutions of the equation. Check your solutions in the original equation. (Enter your answers as a comma-separated list.)
x4 − 8x2 + 7 = 0

Answer by lwsshak3(11628)

Find all solutions of the equation. Check your solutions in the original equation. (Enter your answers as a comma-separated list.)
x4 − 8x2 + 7 = 0
(x^2-7)(x^2-1)=0
..
x^2-7=0
x^2=7
x=±√7
..
x^2-1=0
x^2=1
x=±1
solutions:
x=√7, -√7, 1, -1
..
check:
f(-√7): x^4−8x2+7=49-56+7=0
f(√7): x^4−8x2+7=49-56+7=0
f(-1): x^4−8x2+7=1-8+7=0
f(1): x^4−8x2+7=1-8+7=0